# Integrable measure equivalence rigidity of right-angled Artin groups via quasi-isometry

### v1

@@ Abstract
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$. We prove that if a countable group $H$ with bounded torsion is measure equivalent to $G$, with an $L^{1}$-integrable measure equivalence cocycle towards $G$, then $H$ is finitely generated and quasi-isometric to $G$. In particular, through work of Kleiner and the second-named author, $H$ acts properly and cocompactly on a $\mathrm{CAT}(0)$ cube complex which is quasi-isometric to $G$ and equivariantly projects to the right-angled building of $G$. As a consequence of work of the second-named author, we derive a superrigidity theorem in integrable measure equivalence for an infinite class of right-angled Artin groups, including those whose defining graph is an $n$-gon with $n\geq 5$. In contrast, we also prove that if a right-angled Artin group $G$ with $|\operatorname{Out}(G)|<+\infty$ splits non-trivially as a product, then there does not exist any locally compact group which contains all groups $H$ that are $L^{1}$-measure equivalent to $G$ as lattices, even up to replacing $H$ by a finite-index subgroup and taking the quotient by a finite normal subgroup.
@@ 1.1 Background, history and motivation
By using a notion of uniform measure equivalence to study the large-scale geometry of amenable groups [Sha04], Shalom strengthened the bridge between measure equivalence and quasi-isometry. In the same spirit of imposing a quantitative control on the word length of a cocycle naturally associated to the measure equivalence, Bader–Furman–Sauer coined the notion of integrable measure equivalence, reviewed at the beginning of Section 1.2. For this notion, they established new rigidity theorems for some rank $1$ lattices, including all lattices in $\mathrm{Isom}(\mathbb{H}_{\mathbb{R}}^{n})$ with $n\geq 3$, and all cocompact lattices in $\mathrm{Isom}(\mathbb{H}_{\mathbb{R}}^{2})$ – for the latter rigidity fails for (standard) measure equivalence [BFS13]. The idea behind integrable measure equivalence finds its roots in the work of Margulis [Mar91], where integrability conditions on lattices appear to be crucial in induction arguments, see also [Sha00]. Integrable measure equivalence retains more geometric information about the group, like growth [Aus16b] or the isoperimetric profile [DKLMT22]. On the ergodic side, integrability conditions on orbit equivalence cocycles (closely related to measure equivalence cocycles) already appeared in Belinskaya’s theorem [Bel68] regarding actions of $\mathbb{Z}$, and have also been studied in connection to ergodic notions like entropy [Aus16a, KL21, KL22].
@@ 1.2 Integrable measure equivalence versus quasi-isometry for RAAGs
We say that a group $H$ has bounded torsion if there is a bound on the cardinality of its finite subgroups. Our main theorem is the following.
@@
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, let $H$ be a countable group with bounded torsion. If there exists an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $G$, then $H$ is finitely generated and quasi-isometric to $G$.
As such, the theorem fails if $H$ is allowed to have unbounded torsion. One example comes from infinitely generated non-uniform lattices in the automorphism group of the universal cover of the Salvetti complex of $G$ (see [HH22a, Section 4.2]). Another example comes from graph products over the defining graph of $G$, with vertex groups isomorphic to $\oplus_{\mathbb{N}}\mathbb{Z}/2\mathbb{Z}$ – the odometer gives a measure equivalence (in fact an orbit equivalence) between $\oplus_{\mathbb{N}}\mathbb{Z}/2\mathbb{Z}$ and $\mathbb{Z}$ with an $L^{\infty}$-integrability condition, and the integrability passes to graph products. We do not know any finitely generated examples, however: this is related to the deep and important question regarding the existence of finitely generated non-uniform lattices acting on polyhedral complexes, see [FHT11, Question 33] – we mention that finitely generated non-uniform lattices in products of trees were constructed by Rémy [Ré05] using Kac–Moody groups.
@@ 1.3 Consequences to superrigidity
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, let $H$ be a countable group with bounded torsion. If there exists an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $G$, then $H$ acts properly discontinuously, cocompactly, by cubical automorphisms on a $\mathrm{CAT}(0)$ cube complex which is quasi-isometric to $G$.
@@
We also mention that if the bounded torsion group $H$ is only assumed to be measure equivalent to $G$, with no integrability condition, then we can still derive that $H$ acts cocompactly on a $\mathrm{CAT}(0)$ cube complex (in fact on the right-angled building of $G$), with amenable stabilizers, see Theorem 5.
Under further assumptions on the defining graph of $G$, it is proved in [Hua18, Theorem 1.2], using deep combinatorial insights on special cube complexes by Haglund–Wise [HW12], that all uniform lattices acting on all cube complexes arising in Corollary 2 are virtually special, which is the key for proving that they are commensurable – see also [She22] for recent progress on the relationship between virtual specialness and commensurability. Thus we obtain a superrigidity theorem in integrable measure equivalence for a class of right-angled Artin groups.
@@
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, whose defining graph is star-rigid and does not contain any induced square. Let $H$ be any countable group with bounded torsion. If there exists an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $G$, then $G$ and $H$ are commensurable up to a finite kernel.
@@
The conclusion of Corollary 3 is false whenever the defining graph of $G$ contains an induced square in view of [Hua18, Theorem 1.8]. In this case, the universal cover $X$ of the Salvetti complex of $G$ contains a subcomplex which is a product of two trees. An appropriate irreducible lattice acting on this product of trees can then be extended to a cocompact lattice in $\mathrm{Aut}(X)$ that is not commensurable to $G$ even up to a finite kernel.
@@ 1.4 Lattice envelopes
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, and let $\mathcal{C}_{G}$ be the class of all bounded torsion countable groups $H$ having an $(L^{1},L^{0})$-measure equivalence coupling towards $G$. Motivated by results in Lie groups as discussed before, we ask the following questions. Does there exist a locally compact second countable group $\mathfrak{G}$
@@
It turns out that the answer to the first question is negative whenever $|\operatorname{Out}(G)|<+\infty$, by [HK18, Theorem 6.11] (whose proof constructs a group $H$ that is commensurable to $G$ and has no lattice representation with finite kernel in the same locally compact group as $G$). On the other hand, the second question has a positive answer for the groups appearing in Corollary 3. But our next theorem shows that even the (much more subtle) second question has a negative answer in general, in fact as soon as $G$ splits as a product of two non-cyclic groups. In fact, it even has a negative answer if $\mathcal{C}_{G}$ is replaced with the (possibly smaller) class of all groups $H$ that are strongly commable with $G$, in the sense that there exist finitely generated groups $G=G_{1},\dots,G_{k}=H$ such that for every $i\in\{1,\dots,k-1\}$, the groups $G_{i}$ and $G_{i+1}$ are uniform lattices in a common locally compact second countable group – this definition is a variation over Cornulier’s notion of commability [Cor15].
@@
Let $G$ be the product of two non-cyclic right-angled Artin groups with finite outer automorphism groups (hence $|\operatorname{Out}(G)|<+\infty$).
Then there does not exist any locally compact group $\mathfrak{G}$ such that any torsion-free countable group $H$ which is strongly commable with $G$, has a finite-index subgroup $H^{0}$ with a lattice representation with finite kernel into $\mathfrak{G}$.
@@
Nevertheless, we can still get information on the possible lattice envelopes of a non-cyclic right-angled Artin group $G$ with $|\operatorname{Out}(G)|<+\infty$. More generally, if $H$ is a countable group with bounded torsion which is measure equivalent to $G$, then any lattice embedding of $H$ is cocompact, and in fact every lattice envelope $\mathfrak{H}$ of $H$ is totally disconnected up to a compact kernel (Theorem 10.1). And if there is an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $G$, then more can be said: in this case, there exists a uniformly locally finite $\mathrm{CAT}(0)$ cube complex $Y$ quasi-isometric to $G$ (having the same description as in Corollary 2), and a continuous homomorphism $\mathfrak{H}\to\mathrm{Aut}(Y)$ with compact kernel and cocompact image (Theorem 10.5). We mention that some of these results follow alternatively from the general work of Bader–Furman–Sauer on lattice envelopes [BFS20]. We take a different route, in closer relation to Furman’s ideas for exploiting measure equivalence rigidity towards classification of lattice embeddings [Fur01].
@@ 1.5 Discussion of the proof of the main theorem
Then $H$ acts on $\mathbb{B}$ with amenable vertex stabilizers, and cocompactly provided that $H$ has bounded torsion.
@@
Step 2: From amenable to virtually cyclic stabilizers of rank $1$ vertices. This is the only place where we use our integrability assumption. We use it to show that every vertex of $\mathbb{B}_{\Gamma}$ whose stabilizer for the action of $G_{\Gamma}$ is cyclic, also has a virtually cyclic stabilizer for the action of $H$. For this, using an argument of Escalier and the first-named author [EH], we observe that if the measure equivalence coupling $\Omega$ is $L^{1}$-integrable from $H$ to $G$, then the induced measure equivalence coupling $\Omega_{v}$ between the vertex stabilizers $G_{v},H_{v}$ is also $L^{1}$-integrable from $H_{v}$ to $G_{v}\approx\mathbb{Z}$.
@@
Let $H$ be a countable group with bounded torsion, and assume that there is an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $\mathbb{Z}$.
@@ Acknowledgments.
We thank Uri Bader, Amandine Escalier, Alex Furman, Matthieu Joseph, François Le Maître and Romain Tessera for fruitful related conversations.
The first-named author was funded by the European Union (ERC, Artin-Out-ME-OA, 101040507). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@ 2.1 Right-angled Artin groups
The reader is referred to [Cha07] for an introduction to right-angled Artin groups. Recall that $\Lambda\subset\Gamma$ is an induced subgraph if vertices of $\Lambda$ are adjacent in $\Lambda$ if and only if they are adjacent in $\Gamma$. Each induced subgraph $\Lambda\subset\Gamma$ yields an injective homomorphism $G_{\Lambda}\to G_{\Gamma}$ whose image is called a standard parabolic subgroup of type $\Lambda$. A parabolic subgroup of $G_{\Gamma}$ is a conjugate of some standard parabolic subgroup. A standard coset of type $\Lambda$ is a left coset of the form $gG_{\Lambda}$. A standard abelian subgroup of $G_{\Gamma}$ is a standard parabolic subgroup whose type is a complete subgraph. In particular, the trivial subgroup is a standard abelian subgroup, whose type is the empty set. A standard flat of type $\Lambda$ in $G_{\Gamma}$ is a left coset of a standard abelian subgroup of type $\Lambda$, and its dimension is the rank of this abelian subgroup. One-dimensional standard flats are also called standard lines. Zero-dimensional standard flats are exactly elements in $G_{\Gamma}$. The star of a vertex $v$ in $\Gamma$, denoted by $\mathrm{st}(v)$, is the induced subgraph spanned by $v$ and all its adjacent vertices. Its link $\mathrm{lk}(v)$ is the induced subgraph spanned by all vertices that are adjacent to $v$. The orthogonal $\Lambda^{\perp}$ of an induced subgraph $\Lambda\subseteq\Gamma$ is the induced subgraph of $\Gamma$ spanned by all vertices in $V\Gamma\setminus V\Lambda$ which are adjacent to every vertex of $\Lambda$. For example $\mathrm{lk}(v)=\{v\}^{\perp}$. The following proposition was proved by Charney, Crisp and Vogtmann, building on work of Godelle [God03].
@@
The group $G_{\Gamma}$ acts geometrically, i.e. properly discontinuously and cocompactly, on a $\mathrm{CAT}(0)$ cube complex $X_{\Gamma}$, defined as follows. The 1-skeleton of $X_{\Gamma}$ is the Cayley graph $C_{\Gamma}$ of $G_{\Gamma}$ for its standard generating set $S$. It is equipped with the usual orientation and labeling of edges in a Cayley graph by elements of $S$. We then glue a square to each 4-cycle in the Cayley graph to obtain the 2-skeleton of $X_{\Gamma}$, then attach a $3$-cube to each copy of the boundary of a $3$-cube (with the obvious attaching maps), and more generally by induction on $k$, we glue a $k$-cube on each copy of the boundary of a $k$-cube. This process terminates after finitely many steps and results in a finite-dimensional $\mathrm{CAT}(0)$ cube complex [CD95, Section 3], on which $G_{\Gamma}$ acts geometrically. We will identify $G_{\Gamma}$ with the 0-skeleton of $X_{\Gamma}$. The Salvetti complex of $G_{\Gamma}$, introduced in [Sal87] and denoted $S_{\Gamma}$, is defined to be the quotient $G_{\Gamma}\backslash X_{\Gamma}$ – and $X_{\Gamma}$ is the universal cover of $S_{\Gamma}$. Note that the 2-skeleton of $S_{\Gamma}$ is exactly the presentation complex of $G_{\Gamma}$. Hence we label each edge of $S_{\Gamma}$ by a generator of $G_{\Gamma}$ and orient each edge. Each complete subgraph $\Delta$ of $\Gamma$ with $n$ vertices gives a copy of the $n$-dimensional torus in $S_{\Gamma}$ as a subcomplex, and $S_{\Gamma}$ is a union of these torus subcomplexes.

### v2

@@ Abstract
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$. We prove that if a countable group $H$ with bounded finite subgroups is measure equivalent to $G$, with an $L^{1}$-integrable measure equivalence cocycle towards $G$, then $H$ is finitely generated and quasi-isometric to $G$. In particular, through work of Kleiner and the second-named author, $H$ acts properly and cocompactly on a $\mathrm{CAT}(0)$ cube complex which is quasi-isometric to $G$ and equivariantly projects to the right-angled building of $G$.
As a consequence of work of the second-named author, we derive a superrigidity theorem in integrable measure equivalence for an infinite class of right-angled Artin groups, including those whose defining graph is an $n$-gon with $n\geq 5$. In contrast, we also prove that if a right-angled Artin group $G$ with $|\operatorname{Out}(G)|<+\infty$ splits non-trivially as a product, then there does not exist any locally compact group which contains all groups $H$ that are $L^{1}$-measure equivalent to $G$ as lattices, even up to replacing $H$ by a finite-index subgroup and taking the quotient by a finite normal subgroup.
@@ 1.1 Background, history and motivation
By using a notion of uniform measure equivalence to study the large-scale geometry of amenable groups [Sha04], Shalom strengthened the bridge between measure equivalence and quasi-isometry. In the same spirit of imposing a quantitative control on the word length of a cocycle naturally associated to the measure equivalence, Bader–Furman–Sauer coined the notion of integrable measure equivalence, reviewed at the beginning of Section 1.2. For this notion, they established new rigidity theorems for some rank $1$ lattices, including all lattices in $\mathrm{Isom}(\mathbb{H}_{\mathbb{R}}^{n})$ with $n\geq 3$, and all cocompact lattices in $\mathrm{Isom}(\mathbb{H}_{\mathbb{R}}^{2})$ – for the latter rigidity fails for (standard) measure equivalence [BFS13]. The idea behind integrable measure equivalence finds its roots in the work of Margulis [Mar91], where integrability conditions on lattices appear to be crucial in induction arguments, see also [Sha00]. Integrable measure equivalence retains more geometric information about the group, like growth [Aus16b] or the isoperimetric profile [DKLMT22]. On the ergodic side, integrability conditions on orbit equivalence cocycles (closely related to measure equivalence cocycles) already appeared in Belinskaya’s theorem [Bel68] regarding actions of $\mathbb{Z}$, and have also been studied in connection to ergodic notions like entropy [Aus16a, KL21, KL24].
@@ 1.2 Integrable measure equivalence versus quasi-isometry for RAAGs
We say that a group $H$ has bounded finite subgroups if there is a bound on the cardinality of its finite subgroups. Our main theorem is the following.
@@
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, let $H$ be a countable group with bounded finite subgroups.
If there exists an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $G$, then $H$ is finitely generated and quasi-isometric to $G$.
As such, the theorem fails if $H$ is allowed to have unbounded finite subgroups. This comes from the existence of infinitely generated non-uniform lattices in the automorphism group of the universal cover of the Salvetti complex of $G$, whenever $G$ is a non-abelian right-angled Artin group (see [HH22a, Section 4.2]). Another example of groups $H$ with unbounded finite subgroups having an $(L^{1},L^{0})$-measure equivalence coupling towards $G$ is given by graph products over the defining graph of $G$, with vertex groups isomorphic to $\oplus_{\mathbb{N}}\mathbb{Z}/2\mathbb{Z}$: indeed, the odometer gives a measure equivalence (in fact an orbit equivalence) between $\oplus_{\mathbb{N}}\mathbb{Z}/2\mathbb{Z}$ and $\mathbb{Z}$ with an $L^{\infty}$-integrability condition, and the integrability passes to graph products. We do not know any finitely generated examples, however. This is related to the deep and important question raised in [FHT11, Question 33], asking which polyhedral complexes admit both finitely generated and non-finitely-generated non-uniform lattices.
@@ 1.3 Consequences to superrigidity
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, let $H$ be a countable group with bounded finite subgroups.
If there exists an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $G$, then $H$ acts properly discontinuously, cocompactly, by cubical automorphisms on a $\mathrm{CAT}(0)$ cube complex which is quasi-isometric to $G$.
@@
We also mention that if the group $H$ (with bounded finite subgroups) is only assumed to be measure equivalent to $G$, with no integrability condition, then we can still derive that $H$ acts cocompactly on a $\mathrm{CAT}(0)$ cube complex (in fact on the right-angled building of $G$), with amenable stabilizers, see Theorem 5.
Under further assumptions on the defining graph of $G$, it is proved in [Hua18, Theorem 1.2], using deep combinatorial insights on special cube complexes by Haglund–Wise [HW12], that all uniform lattices acting on all cube complexes arising in Corollary 2 are virtually special, which is the key for proving that they are commensurable – see also [She24] for recent progress on the relationship between virtual specialness and commensurability. Thus we obtain a superrigidity theorem in integrable measure equivalence for a class of right-angled Artin groups.
@@
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, whose defining graph is star-rigid and does not contain any induced square. Let $H$ be any countable group with bounded finite subgroups.
If there exists an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $G$, then $G$ and $H$ are commensurable up to a finite kernel.
@@
The conclusion of Corollary 3 is false whenever the defining graph of $G$ contains an induced square in view of [Hua18, Theorem 1.8] (regardless of whether $\operatorname{Out}(G)$ is finite or not). In this case, the universal cover $X$ of the Salvetti complex of $G$ contains a subcomplex which is a product of two trees. An appropriate irreducible lattice acting on this product of trees can then be extended to a cocompact lattice in $\mathrm{Aut}(X)$ that is not commensurable to $G$ even up to a finite kernel.
@@ 1.4 Lattice envelopes
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, and let $\mathcal{C}_{G}$ be the class of all countable groups $H$ with bounded finite subgroups having an $(L^{1},L^{0})$-measure equivalence coupling towards $G$. Motivated by results in Lie groups as discussed before, we ask the following questions. Does there exist a locally compact second countable group $\mathfrak{G}$
@@
It turns out that the answer to the first question is negative whenever $|\operatorname{Out}(G)|<+\infty$, by [HK18, Theorem 6.11] (whose proof constructs a group $H$ that is commensurable to $G$ and has no lattice representation with finite kernel in the same locally compact group as $G$). On the other hand, the second question has a positive answer for the groups appearing in Corollary 3. But our next theorem provides a class of right-angled Artin groups $G$ where even the second question as a negative answer, namely: the answer to the second question is negative whenever $|\operatorname{Out}(G)|<+\infty$ and $G$ splits as a product of two non-cyclic groups. In fact, for these groups $G$, the second question even has a negative answer if $\mathcal{C}_{G}$ is replaced with the (possibly smaller) class of all groups $H$ that are strongly commable with $G$, in the sense that there exist finitely generated groups $G=G_{1},\dots,G_{k}=H$ such that for every $i\in\{1,\dots,k-1\}$, the groups $G_{i}$ and $G_{i+1}$ are uniform lattices in a common locally compact second countable group – this definition is a variation over Cornulier’s notion of commability [Cor15].
@@
Let $G$ be the direct product of two non-cyclic right-angled Artin groups with finite outer automorphism groups (hence $|\operatorname{Out}(G)|<+\infty$).
Then there does not exist any locally compact group $\mathfrak{G}$ such that any torsion-free countable group $H$ which is strongly commable with $G$, has a finite-index subgroup $H^{0}$ with a lattice embedding into $\mathfrak{G}$.
To prove Theorem 4, we construct an infinite sequence of torsion-free countable groups $H_{n}$ that are all strongly commable to $G$, and which cannot all virtually embed as lattices in a common locally compact group. We do not know, however, whether we can find two torsion-free countable groups $H_{1},H_{2}$, both commable with $G$, such that no finite-index subgroups $H_{1}^{0}$ of $H_{1}$ and $H_{2}^{0}$ of $H_{2}$ have a lattice embedding in a common locally compact group.
@@
Nevertheless, we can still get information on the possible lattice envelopes of a non-cyclic right-angled Artin group $G$ with $|\operatorname{Out}(G)|<+\infty$. More generally, if $H$ is a countable group with bounded finite subgroups which is measure equivalent to $G$, then any lattice embedding of $H$ is cocompact, and in fact every lattice envelope $\mathfrak{H}$ of $H$ is totally disconnected up to a compact kernel (Theorem 10.1). And if there is an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $G$, then more can be said: in this case, there exists a uniformly locally finite $\mathrm{CAT}(0)$ cube complex $Y$ quasi-isometric to $G$ (having the same description as in Corollary 2), and a continuous homomorphism $\mathfrak{H}\to\mathrm{Aut}(Y)$ with compact kernel and cocompact image (Theorem 10.5). We mention that some of these results follow alternatively from the general work of Bader–Furman–Sauer on lattice envelopes [BFS20]. We take a different route, in closer relation to Furman’s ideas for exploiting measure equivalence rigidity towards classification of lattice embeddings [Fur01]. On the topic of lattice envelopes of right-angled Artin groups, we also refer the reader to the recent work of Caprace–de Medts [CdM24].
@@ 1.5 Discussion of the proof of the main theorem
Then $H$ acts on $\mathbb{B}$ with amenable vertex stabilizers, and cocompactly provided that $H$ has bounded finite subgroups.
@@
Step 2: From amenable to virtually cyclic stabilizers of rank $1$ vertices. This is the only place where we use our integrability assumption. We use it to show that every vertex of $\mathbb{B}_{\Gamma}$ whose stabilizer for the action of $G_{\Gamma}$ is cyclic, also has a virtually cyclic stabilizer for the action of $H$. For this, using an argument of Escalier and the first-named author [EH24], we observe that if the measure equivalence coupling $\Omega$ is $L^{1}$-integrable from $H$ to $G$, then the induced measure equivalence coupling $\Omega_{v}$ between the vertex stabilizers $G_{v},H_{v}$ is also $L^{1}$-integrable from $H_{v}$ to $G_{v}\approx\mathbb{Z}$.
@@
Let $H$ be a countable group with bounded finite subgroups, and assume that there is an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $\mathbb{Z}$.
@@ Acknowledgments.
We thank Uri Bader, Pierre-Emmanuel Caprace, Amandine Escalier, Alex Furman, Matthieu Joseph, François Le Maître and Romain Tessera for fruitful related conversations. We are also grateful to Yves Cornulier for helpful comments on an earlier version of this paper. Finally, we thank the referee for their helpful comments on our paper.
The first-named author was funded by the European Union (ERC, Artin-Out-ME-OA, 101040507). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. A CC-BY public copyright license has been applied by the authors to the present document and will be applied to all subsequent versions arising from this submission, in accordance with the grant’s open access conditions.
@@ 2.1 Right-angled Artin groups
The reader is referred to [Cha07] for an introduction to right-angled Artin groups.
Recall that $\Lambda\subset\Gamma$ is an induced subgraph if vertices of $\Lambda$ are adjacent in $\Lambda$ if and only if they are adjacent in $\Gamma$. Each induced subgraph $\Lambda\subset\Gamma$ yields an injective homomorphism $G_{\Lambda}\to G_{\Gamma}$ whose image is called a standard parabolic subgroup of type $\Lambda$. A parabolic subgroup of $G_{\Gamma}$ is a conjugate of some standard parabolic subgroup. A standard coset of type $\Lambda$ is a left coset of the form $gG_{\Lambda}$. A standard abelian subgroup of $G_{\Gamma}$ is a standard parabolic subgroup whose type is a complete subgraph. In particular, the trivial subgroup is a standard abelian subgroup, whose type is the empty set. A standard flat of type $\Lambda$ in $G_{\Gamma}$ is a left coset of a standard abelian subgroup of type $\Lambda$, and its dimension is the rank of this abelian subgroup. One-dimensional standard flats are also called standard lines. Zero-dimensional standard flats are exactly elements in $G_{\Gamma}$.
The star of a vertex $v$ in $\Gamma$, denoted by $\mathrm{st}(v)$, is the induced subgraph spanned by $v$ and all its adjacent vertices. Its link $\mathrm{lk}(v)$ is the induced subgraph spanned by all vertices that are adjacent to $v$.
The orthogonal $\Lambda^{\perp}$ of an induced subgraph $\Lambda\subseteq\Gamma$ is the induced subgraph of $\Gamma$ spanned by all vertices in $V\Gamma\setminus V\Lambda$ which are adjacent to every vertex of $\Lambda$. For example $\mathrm{lk}(v)=\{v\}^{\perp}$. The following proposition was proved by Charney, Crisp and Vogtmann, building on work of Godelle [God03].
@@
[…]